i need help....choose a method to slove the linear system. ( the two equations are together)
8x+y=15
2x+2y=9

Found 2 solutions by elima, Paul:
Answer by elima(1433)

8x+y=15
2x+2y=9
There are different ways to solve this,I'll solve by substitution.
Solve eq.1 8x+y=15 for y;
8x+y-8x=15-8x
y=15-8x
Now substitute this for y in the 2nd eq.;
2x+2(15-8x)=9
2x+30-16x=9
-14x+30=9
-14x=9-30
-14x=-21
x=1.5
Now that we have x we can plug that into eq.1 to get y;
8(1.5)+y=15
12+y=15
y=15-12
y=3
So the answer is;
x=1.5
y=3
Hope you understand!
=)

Answer by Paul(988)

You can either solve it with elimination method or subsitution method
Subsitution method
y=15-8x
2x+2y=9 sub in for y
2x+2(15-8x)=9
2x+30-16x=9
-14x=-21
x=1.5


Elimination method:
8x+y=15 (1)
2x+2y=9 (2) ---> multiply (1) by 2
16x+2y=30
2x+2y=9 ---> nopw subbtract 1 by 2

14x=21
x=1.5
Now just solve for y.